OpenStudy (anonymous):

What is the slope of the line that passes through (3,7) and (-1,1)

OpenStudy (amistre64):

move the points so that one of them is at the origin (3,7) and (-1,1) 1 -1 1 -1 ------------- 4 6 0 0 then the slope is defined by the none zero point as: y/x

OpenStudy (amistre64):

@terenzreignz

terenzreignz (terenzreignz):

Point is well taken :)

terenzreignz (terenzreignz):

Both points XD

OpenStudy (anonymous):

Since line passes through (3,7) and (-1,1) therefore \[x_{1}=3 \] \[x_{2}=-1\] \[y_{1}=7\] \[y_{2}=1\] Slope of line is given as \[slope =\frac{y_{2}-y_{1}}{x_{2}-x_{1}}=\frac{1-7}{-1-3}=\frac{-6}{-4}=\frac{3}{2}\]